OpenStudy (zyi6):

Harry has a jar that contains 75% of apple juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of apple juice and water in the jar. Define the variable used in the equation. (5 points) Part B: How many liters of apple juice are present in the jar? Show your work. (5 points)

OpenStudy (zyi6):

@tkhunny , @texaschic101 @TheSmartOne

OpenStudy (tkhunny):

Nice problem. Please show your work and share your ideas. What quiz is this? Should you be getting outside help?

OpenStudy (zyi6):

This is off of a practice test. All i know is this: 75% = 3 / 4 1 liter = 1 / 4 togehter, you would get 4 liters. how do i set up the eqauation though?

OpenStudy (tkhunny):

75% Apple Juice 25% Water = 1 liter water Always, always, ALWAYS define, write down, and remember useful and necessary variables. How many liters in total? No idea. Name it. Define it. Describe it. T = Total Number of Liters of stuff. Once that has been done, can we now talk about equations? Rule #1 - Show your work up front. Don't make people guess what you are doing. Show them!

OpenStudy (zyi6):

noted. now that i have provided you with my information that I know, can you walk me through this?

OpenStudy (tkhunny):

I just did. T = Total Number of Liters of stuff. Now, go ahead and write an equation expressing the amount of Apple Juice. Water: 0.25T = 1 liter

OpenStudy (zyi6):

T = 0.25t + 0.75t???????